OpenStudy (anonymous):

A rectangle has area 49 m2. (a) Express the perimeter of the rectangle as a function of the length (L) of one of its sides.

OpenStudy (anonymous):

(b) State its domain.

OpenStudy (anonymous):

i got 2x + 32/x , but don't know if its right nor how to find the domain

OpenStudy (anonymous):

A=L*W=49 P=2(L+W)

OpenStudy (anonymous):

W=49/L P=2(L+49/L)

OpenStudy (anonymous):

do you know how to find the domain

OpenStudy (anonymous):

domain is the set of those values of variable for which the function is always defined

OpenStudy (anonymous):

so how would it be for this equation